a proton moving horizontally at 3.7 x 10 to the fifth power ms and thirst and a uniform magnetic field of 0.08 e60 directed upward at 35 degrees with the horizontal find the magnitude of the magnetic force acting on the proton

1
Expert's answer
2022-05-22T14:38:40-0400
"F=qvB\sin\theta\\
=1.6*10^{-19}*3.7*10^5*0.08*\sin 35^\circ\\
=2.7*10^{-15}\:\rm N"